OpenStudy (anonymous):

Find the derivative of y=(x^2 +4)^5 HELP PLEASE I WILL AWARD MEDALS!

OpenStudy (tkhunny):

Have you considered the Chain Rule?

OpenStudy (anonymous):

Yes, can you help me get started please

OpenStudy (anonymous):

@satellite73 @Loser66

OpenStudy (anonymous):

You have something raised to the fifth power. You know how to find the derivative of something raised to the fifth power. Then you have to apply the chain rule, multiplying by the derivative of that "something."

OpenStudy (anonymous):

So would I just get the expanded form and then multiply it out?

OpenStudy (anonymous):

10x^9+160x^7+960x^5 +2560X^3+2560 x

OpenStudy (anonymous):

@creeksider

OpenStudy (anonymous):

y=(x^2 +4)^5 y' = 5(2x)(x^2+4)^4

OpenStudy (anonymous):

the derivative of "something" to the power of 5 is \(5\text{something}^4\times\text{derivative of something}\)

OpenStudy (anonymous):

We can give you the answer but that won't help if you don't understand how to get it. You need to go back over the chain rule, because it's an essential tool in calculus and when you understand it, problems like this are a snap. satellite73 gave you an outline of how it works for this problem.

OpenStudy (nali):

if you need help understanding the chain rule, here is a helpful video: https://www.khanacademy.org/math/calculus/differential-calculus/chain_rule/v/the-chain-rule

OpenStudy (anonymous):

(dy(x))/(dx) = 10 x (4+x^2)^4